OpenStudy (anonymous):

(3+√2ˏ6+√2 ) (3−√2ˏ6−√2 ) = Find the distance between two points ? :)

OpenStudy (astrophysics):

You can use the distance formula \[\huge d = \sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] where your coordinates represent \[(x_1, y_1) ~~~\text{and}~~~(x_2,y_2)\] respectively.

OpenStudy (anonymous):

@Loonie here is your answer

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Best of Luck

OpenStudy (anonymous):

OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution. Have homework problems? We can help you - as long as you are looking to learn how to get your solution (and not just the answer). Please do not copy and paste your problems. If people are going to spend time to help you, you should at least spend the time to ask your question as clearly as possible.

OpenStudy (anonymous):

why ? this is the answer ? d = √(8+8) = √16 = 4 ?

OpenStudy (usukidoll):

@Loonie this isn't aimed at you.... It's aimed at @geniushr What @peachpi wrote... That is from the Openstudy Code of Conduct which was obviously broken thanks to the direct answer post. This website is solely used to guide users through the process of learning the material.

OpenStudy (usukidoll):

Stop giving direct answers! I bet you're trying to reach 80 SS fast because of what I just typed in OpenStudy Feedback earlier. Isn't it?

OpenStudy (usukidoll):

right @peachpi ?

OpenStudy (anonymous):

yes I probably should have tagged @geniushr to make that clear

OpenStudy (anonymous):

d = √{[(3−√2)-(3+√2)]² + [(6−√2)-(6+√2)]²} d = √[(-2√2)²+(-2√2)²]

OpenStudy (anonymous):

I know it :) Thanks @geniushr

OpenStudy (anonymous):

Best of Luck dear :)

OpenStudy (michele_laino):

@geniushr Please note that you can not give to the asker the direct answer, since you have to respect the Code Of Conduct

OpenStudy (anonymous):

@Michele_Laino Sorry I was not fimilier of it...

OpenStudy (usukidoll):

http://openstudy.com/code-of-conduct

OpenStudy (michele_laino):

@geniushr ok!, please give your explanation to that question

OpenStudy (anonymous):

@Michele_Laino she already gave the explanation :)

OpenStudy (michele_laino):

I meant, that we have to guide the asker in understanding of his/her question @Loonie

OpenStudy (anonymous):

I know how to solve distance formula with square root :)

OpenStudy (anonymous):

@Michele_Laino ok :)))

OpenStudy (michele_laino):

It is ok if @geniushr will follow the Code of Conduct :)

OpenStudy (michele_laino):

from now onwards :)

OpenStudy (anonymous):

Yes dear I shall follow the code of conduct afterwards

OpenStudy (michele_laino):

thanks! :) @geniushr

OpenStudy (anonymous):

thanks :)